OpenStudy (anonymous):

Use the distributive property to rewrite the expression and solve. 9 * 72 = 9 * (Blank space) + 9 * 2 = (Another blank space)

OpenStudy (anonymous):

I need to fill in both blank spaces, and my math book doesn't really clarify how

OpenStudy (anonymous):

\[9\times 72=9\times ?+9\times 2\]

OpenStudy (anonymous):

\[72=2\times ?\]i.e what is half of \(72\)?

OpenStudy (anonymous):

Oh my god sorry, my computer jammed. Half of 72 is 36. @satellite73

OpenStudy (anonymous):

ok that is the answer to the fill in the blank question

OpenStudy (anonymous):

Wait, there are two blanks, they're both 36?

OpenStudy (anonymous):

no, the last blank space is for whatever \(9\times 72\) is

OpenStudy (anonymous):

that's 648

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Alright. I have just one more question, can you help me with it? @satellite73